# Efficient Fourier Filtering Network with Contrastive Learning for UAV-based Unaligned Bi-modal Salient Object Detection

### v2

@@ Abstract
Unmanned aerial vehicle (UAV)-based bi-modal salient object detection (BSOD) aims to segment salient objects in a scene utilizing complementary cues in unaligned RGB and thermal image pairs. However, the high computational expense of existing UAV-based BSOD models limits their applicability to real-world UAV devices. To address this problem, we propose an efficient Fourier filter network with contrastive learning that achieves both real-time and accurate performance. Specifically, we first design a semantic contrastive alignment loss to align the two modalities at the semantic level, which facilitates mutual refinement in a parameter-free way. Second, inspired by the fast Fourier transform that obtains global relevance in linear complexity, we propose synchronized alignment fusion, which aligns and fuses bi-modal features in the channel and spatial dimensions by a hierarchical filtering mechanism. Our proposed model, AlignSal, reduces the number of parameters by 70.0%, decreases the floating point operations by 49.4%, and increases the inference speed by 152.5% compared to the cutting-edge BSOD model (i.e., MROS). Extensive experiments on the UAV RGB-T 2400 and seven bi-modal dense prediction datasets demonstrate that AlignSal achieves both real-time inference speed and better performance and generalizability compared to nineteen state-of-the-art models across most evaluation metrics. In addition, our ablation studies further verify AlignSal’s potential in boosting the performance of existing aligned BSOD models on UAV-based unaligned data. The code is available at: https://github.com/JoshuaLPF/AlignSal.
Index Terms:
Bi-modal salient object detection, contrastive alignment, fast Fourier transform, modal alignment, unmanned aerial vehicles
@@ I Introduction
Figure 1: Examples from UAV RGB-T 2400 dataset [1]. RGB images typically have higher resolution and a wider field of view compared to thermal images.
@@
Unmanned aerial vehicle (UAV)-based object detection methods have been broadly applied in various fields, such as emergency rescue [2, 3], plant maintenance [4], and electrical overhaul [5]. Despite the impressive achievement of existing detection models [2, 6, 7, 8], methods that rely solely on RGB [2, 6] or thermal imaging [7, 8] face challenges in scenes with low illumination, occlusion, and background clutter. The two modalities have their own characteristics. While the RGB modality provides rich color and texture information about a scene, the thermal modality utilizes temperature differences to identify objects independently of lighting conditions. Therefore, recent works [9, 1] have attempted to combine the strengths of RGB and thermal modalities to improve UAV-based object detection in complex scenes.
@@
In practice, the resolution and scale of images captured by the RGB and thermal cameras mounted on UAVs may differ, resulting in unaligned bi-modal images, as shown in Fig. 1. These differences lead to varying proportions and spatial locations of the same object between the two modalities. However, most existing bi-modal object detection datasets [10, 11, 12] are manually aligned, avoiding many real-world challenges due to the misalignment between the two modalities. Consequently, models designed based on these aligned datasets [13, 14, 15, 16, 17, 18, 12, 19, 20, 21, 22, 23] usually underperform when applied to unaligned data under realistic conditions.
@@
To address these issues, Song et al. [1] constructed the first unaligned bi-modal salient object detection (BSOD) dataset, UAV RGB-T 2400, based on the UAV view and proposed a paradigm model named MROS. MROS achieves remarkable performance by aligning the two modalities, then performing modal fusion, followed by decoding and generating predictions. However, two primary challenges remain. First, UAVs are often used for real-time scene detection, but the complexity of MROS - with 90.4M parameters and 45.3G floating point operations - places a significant burden on the UAV’s processor and memory. In particular, modal alignment and fusion modules have high computational demands, limiting the overall inference speed. Second, since UAVs operate at high altitudes, the scale of the objects captured is small, which increases spatial location offsets of small-scale objects between the two modalities. The alignment strategies in MROS, based on convolutional attention operations [24], struggle to handle large spatial offsets effectively due to their limited receptive fields.
@@
In this paper, we propose a two-step solution to address the aforementioned limitations of the current BSOD methods. First, we aim to develop parameter-free learning strategies that reduce model complexity while achieving essential modal alignment. Second, we simultaneously perform weak bi-modal alignment and fusion, further minimizing testing complexity to improve inference speed. By integrating bi-modal features into a shared feature space, we aim to establish a multi-dimensional, multi-filtering mechanism based on the mutual relationships between the two modalities on a global scale to address varying offset scales and leverage complementary bi-modal information effectively. To achieve this, in this work, we propose AlignSal, an efficient Fourier filtering network with contrastive learning for accurate and real-time UAV-based unaligned BSOD.
@@
Specifically, AlignSal is composed of two major novelties that contribute to its superior performance. First, we propose the semantic contrastive alignment loss (SCAL) to align RGB and thermal modalities at the semantic level. Inspired by contrastive learning [25, 26, 27, 28, 29], SCAL forces similar local features in RGB and thermal modalities to stay close in the embedding space while pushing the dissimilar features apart. This knowledge exchange process allows both modalities to refine each other, achieving modal alignment while improving their individual representations. Notably, SCAL does not increase computational loads during inference. Second, we propose the synchronized alignment fusion (SAF) module, which leverages the fast Fourier transform to align features in the channel and spatial dimensions and facilitate bi-modal fusion. By modeling channel dependencies, SAF employs a multi-group global filtering mechanism that hierarchically captures spatial offsets and bi-modal salient cues at multiple scales, further aligning and integrating them into the fused feature. The low complexity of SAF ensures real-time performance during inference. With the proposed SCAL and SAF, AlignSal effectively aligns RGB and thermal modalities by leveraging their correlations and complementary strengths, producing accurate saliency maps from the UAV perspective.
@@
  • We propose an efficient model, AlignSal, for UAV-based unaligned BSOD. With detailed ablation studies, we validate the effectiveness of the proposed key components, namely SCAL and SAF. We further demonstrate SCAL’s ability to enhance the performance of existing aligned BSOD models on the unaligned bi-modal data.
  • With extensive experiments conducted on the UAV RGB-T 2400 and seven additional bi-modal dense prediction datasets, we show that our proposed AlignSal achieves better performance and generalizability compared to nineteen existing state-of-the-art models, while achieving real-time inference speed.
@@ II Related Work
II-A Bi-modal Salient Object Detection (BSOD)
@@
With advancements in sensor technology, bi-modal salient object detection (BSOD) has emerged as a prominent research area [30, 31, 32, 33, 34, 35]. The introduction of thermal information enhances the model’s ability to distinguish objects in complex environments [36, 37, 38]. Thermal images capture temperature distributions, highlighting object contours, while RGB images deliver abundant color and contextual information that enables a more intuitive understanding of the scene. To leverage inter-modal complementary cues, existing BSOD models can be categorized into early, middle, and late fusion approaches.
@@
Early fusion models typically aggregate bi-modal image pairs linearly as input [39, 17, 40]. While this strategy reduces computational load, it fails to prevent the inter-modal noise from interfering with each other. In contrast, late fusion models often employ dual- [41, 13] or triple-stream branches [42, 43] that focus on semantic features or decoder feature fusion. However, these approaches overlook features at early levels that retain important fine-grained spatial information. Therefore, the balanced middle fusion strategy has become the dominant approach in BSOD [44, 45, 14, 15, 18, 12, 46, 22, 23]. For instance, Huo et al. [14] proposed an ultra- lightweight network with cross-modal filtering and step-wise refinement, while a global illumination module was introduced [18] to determine the use of the thermal modality. Zhou et al. [22] developed a wavelet-based model that uses knowledge distillation to transfer information from a complex model. Wang et al. [23] proposed an adaptive fusion bank, which utilizes basic fusion schemes to address different challenges simultaneously. Despite the significant advances in these models, their performance is constrained by the limited view of convolution operations. Inspired by the attention mechanism employed by the Transformer [47] which is good at capturing global information, researchers have applied Transformer-based solutions to BSOD [48, 49, 50, 51, 16, 19, 52, 20, 21]. Specifically, recent works [50, 16, 19, 20] employed Transformer-based backbones to model initial multi-level features. Zhou et al. [51] designed an auxiliary task to optimize pixel classification and proposed a Transformer-based decoder. Pang et al. [21] proposed view-mixed attention for integrating cross-modal and cross-level features. Cong et al. [52] introduced position constraints to enhance cross-modal long-range modeling. Although these models effectively address BSOD challenges, they are designed for manually aligned bi-modal image pairs and struggle with unaligned instantaneous data. In contrast, our proposed model, AlignSal, is developed for unaligned BSOD data, which can process unaligned RGB and thermal image pairs in real time and exploit the inter-modal complementarity to accurately identify salient objects.
@@
Recently, Tu et al. [53] simulated weakly aligned data by applying affine transformations to aligned bi-modal data and used dynamic convolution for weakly aligned BSOD. However, such artificially unaligned data does not fully capture real-world conditions. To address this, Song et al. [1] constructed an unaligned BSOD dataset based on UAV views of urban scenes and proposed the MROS model, which performs image alignment followed by fusion. While MROS demonstrates excellent performance, its high computational demands hinder real-time detection capabilities. In order to prioritize both accuracy and real-time performance, we allocate resource-intensive bi-modal alignment to the training phase, which shifts part of the computational burden from inference to training. We also design a hierarchical global filtering mechanism with low complexity to handle modal fusion and alignment. Our model achieves competitive performance compared to state-of-the-art models while enabling real-time detection with manageable complexity.
@@ II-B Contrastive Learning (CL)
Contrastive learning (CL) focuses on learning data representations by analyzing the similarities and differences between samples. In CL, the distance between features of similar samples in the embedding space is minimized while that between dissimilar samples is maximized, facilitating the acquisition of meaningful feature representations [54]. CL has been successfully applied in various fields of computer vision [25, 55, 28, 56, 49, 29]. For instance, Radford et al. [25] proposed a model for aligning visual and textual information through CL, and Jiang et al. [28] combined visual and textual prompts within a single model via CL. Song et al. [29] aligned RGB and radar features using CL to improve the robustness of fusion. Beyond its use for aligning heterogeneous modalities, CL has also been widely studied to enhance inter-pixel discrimination. Wang et al. [55] extended pixel-level CL from a single image to neighboring images. Sung et al. [56] proposed contextual-based CL to obtain clear feature distinctions. Wu et al. [49] utilized CL to increase the separation between foreground and background pixels in predictions. In this paper, to address the misalignment of RGB and thermal modalities, we leverage CL to align bi-modal features at the semantic level and address UAV-based unaligned BSOD.
Figure 2: The overall framework of the proposed AlignSal. The RGB and thermal images are first fed to a dual-stream encoder, which extracts initial bi-modal features $\left\{{f_{i}^{r}}\right\}_{i=1}^{4}$ and $\left\{{f_{i}^{t}}\right\}_{i=1}^{4}$. During training, the semantic contrastive alignment loss (SCAL) facilitates the exchange of information between the RGB semantic feature $f_{4}^{r}$ and the thermal semantic feature $f_{4}^{t}$, essentially achieving bi-modal alignment. The bi-modal features are then aligned and fused at the pixel level by the synchronized alignment fusion (SAF). Finally, a simple decoder integrates the fused features $\left\{{f_{i}^{s}}\right\}_{i=1}^{4}$ from high to low levels, generating the decoder features $\left\{{f_{i}^{d}}\right\}_{i=1}^{4}$ and the final prediction $\mathcal{S}$.
@@ III-A Overview
As shown in Fig. 2, the proposed AlignSal consists of four main components: the dual-stream encoder, the semantic contrastive alignment loss (SCAL), the synchronized alignment fusion (SAF), and the decoder. First, considering that the fast Fourier transform (FFT) offers low complexity while capturing global dependencies, we use the lightweight FFT-based CDFFormer-S18 [57] as the encoder. Leveraging the lower inductive bias introduced by FFT [58, 59, 57, 22], we employ a Siamese architecture [60, 22] for the encoder to extract four levels of RGB features $\left\{{f_{i}^{r}}\right\}_{i=1}^{4}$ and thermal features $\left\{{f_{i}^{t}}\right\}_{i=1}^{4}$, further reducing complexity. During the training phase, the high-level bi-modal features $f_{4}^{r}$ and $f_{4}^{t}$, rich in semantic information, are aligned by SCAL to achieve a consistent semantic distribution. Simultaneously, these bi-modal features are fed into the corresponding SAF in pairs to achieve pixel-level alignment while capturing inter-modal complementary information. Finally, the outputs from the SAFs $\left\{{f_{i}^{s}}\right\}_{i=1}^{4}$ are passed through a simple decoder to progressively generate up-sampling features $\left\{{f_{i}^{d}}\right\}_{i=1}^{4}$ and the saliency map $\mathcal{S}$.

### v3

# Efficient Fourier Filtering Network with Contrastive Learning for AAV-based Unaligned Bimodal Salient Object Detection

@@ Abstract
Autonomous aerial vehicle (AAV)-based bimodal salient object detection (BSOD) aims to segment salient objects in a scene utilizing complementary cues in unaligned RGB and thermal image pairs. However, the high computational expense of existing AAV-based BSOD models limits their applicability to real-world AAV devices. To address this problem, we propose an efficient Fourier filter network with contrastive learning that achieves both real-time and accurate performance. Specifically, we first design a semantic contrastive alignment loss to align the two modalities at the semantic level, which facilitates mutual refinement in a parameter-free way. Second, inspired by the fast Fourier transform that obtains global relevance in linear complexity, we propose synchronized alignment fusion, which aligns and fuses bimodal features in the channel and spatial dimensions by a hierarchical filtering mechanism. Our proposed model, AlignSal, reduces the number of parameters by 70.0%, decreases the floating point operations by 49.4%, and increases the inference speed by 152.5% compared to the cutting-edge BSOD model (i.e., MROS). Extensive experiments on the AAV RGB-T 2400 and seven bimodal dense prediction datasets demonstrate that AlignSal achieves both real-time inference speed and better performance and generalizability compared to nineteen state-of-the-art models across most evaluation metrics. In addition, our ablation studies further verify AlignSal’s potential in boosting the performance of existing aligned BSOD models on AAV-based unaligned data. The code is available at: https://github.com/JoshuaLPF/AlignSal.
@@ I Introduction
Figure 1: Examples from AAV RGB-T 2400 dataset [10315195]. RGB images typically have higher resolution and a wider field of view compared to thermal images.
@@
Autonomous aerial vehicle (AAV)-based object detection methods have been broadly applied in various fields, such as emergency rescue [bo2022basnet, xu2024yoloow], plant maintenance [hu2023detection], and electrical overhaul [senthilnath2021bs]. Despite the impressive achievement of existing detection models [bo2022basnet, zhang2023cfanet, li2020comnet, fang2024scinet], methods that rely solely on RGB [bo2022basnet, zhang2023cfanet] or thermal imaging [li2020comnet, fang2024scinet] face challenges in scenes with low illumination, occlusion, and background clutter. The two modalities have their own characteristics. While the RGB modality provides rich color and texture information about a scene, the thermal modality utilizes temperature differences to identify objects independently of lighting conditions. Therefore, recent works [zhang2022visible, 10315195] have attempted to combine the strengths of RGB and thermal modalities to improve AAV-based object detection in complex scenes.
@@
In practice, the resolution and scale of images captured by the RGB and thermal cameras mounted on AAVs may differ, resulting in unaligned bimodal images, as shown in Fig. 1. These differences lead to varying proportions and spatial locations of the same object between the two modalities. However, most existing bimodal object detection datasets [peng2014rgbd, 9767629, 10003255] are manually aligned, avoiding many real-world challenges due to the misalignment between the two modalities. Consequently, models designed based on these aligned datasets [9454273, 9505635, jin2022moadnet, 9611276, 9803225, 9926193, 10003255, sun2023catnet, 9869666, 10015667, 10127616, wang2024learning] usually underperform when applied to unaligned data under realistic conditions.
@@
To address these issues, Song et al. [10315195] constructed the first unaligned bimodal salient object detection (BSOD) dataset, AAV RGB-T 2400, based on the AAV view and proposed a paradigm model named MROS. MROS achieves remarkable performance by aligning the two modalities, then performing modal fusion, followed by decoding and generating predictions. However, two primary challenges remain. First, AAVs are often used for real-time scene detection, but the complexity of MROS - with 90.4M parameters and 45.3G floating point operations - places a significant burden on the AAV’s processor and memory. In particular, modal alignment and fusion modules have high computational demands, limiting the overall inference speed. Second, since AAVs operate at high altitudes, the scale of the objects captured is small, which increases spatial location offsets of small-scale objects between the two modalities. The alignment strategies in MROS, based on convolutional attention operations [woo2018cbam], struggle to handle large spatial offsets effectively due to their limited receptive fields.
@@
In this paper, we propose a two-step solution to address the aforementioned limitations of the current BSOD methods. First, we aim to develop parameter-free learning strategies that reduce model complexity while achieving essential modal alignment. Second, we simultaneously perform weak bimodal alignment and fusion, further minimizing testing complexity to improve inference speed. By integrating bimodal features into a shared feature space, we aim to establish a multi-dimensional, multi-filtering mechanism based on the mutual relationships between the two modalities on a global scale to address varying offset scales and leverage complementary bimodal information effectively. To achieve this, in this work, we propose AlignSal, an efficient Fourier filtering network with contrastive learning for accurate and real-time AAV-based unaligned BSOD.
@@
Specifically, AlignSal is composed of two major novelties that contribute to its superior performance. First, we propose the semantic contrastive alignment loss (SCAL) to align RGB and thermal modalities at the semantic level. Inspired by contrastive learning [radford2021learning, yang2021taco, he2023align, jiang2024t, song2024contrastalign], SCAL forces similar local features in RGB and thermal modalities to stay close in the embedding space while pushing the dissimilar features apart. This knowledge exchange process allows both modalities to refine each other, achieving modal alignment while improving their individual representations. Notably, SCAL does not increase computational loads during inference. Second, we propose the synchronized alignment fusion (SAF) module, which leverages the fast Fourier transform to align features in the channel and spatial dimensions and facilitate bimodal fusion. By modeling channel dependencies, SAF employs a multi-group global filtering mechanism that hierarchically captures spatial offsets and bimodal salient cues at multiple scales, further aligning and integrating them into the fused feature. The low complexity of SAF ensures real-time performance during inference. With the proposed SCAL and SAF, AlignSal effectively aligns RGB and thermal modalities by leveraging their correlations and complementary strengths, producing accurate saliency maps from the AAV perspective.
@@
  • We propose an efficient model, AlignSal, for AAV-based unaligned BSOD. With detailed ablation studies, we validate the effectiveness of the proposed key components, namely SCAL and SAF. We further demonstrate SCAL’s ability to enhance the performance of existing aligned BSOD models on the unaligned bimodal data.
  • With extensive experiments conducted on the AAV RGB-T 2400 and seven additional bimodal dense prediction datasets, we show that our proposed AlignSal achieves better performance and generalizability compared to nineteen existing state-of-the-art models, while achieving real-time inference speed.
@@ II Related Work
II-A Bimodal Salient Object Detection (BSOD)
@@
With advancements in sensor technology, bimodal salient object detection (BSOD) has emerged as a prominent research area [wang2021cgfnet, gao2021unified, zhao2022self, zhang2021deep, lee2022spsn, wu2023hidanet]. The introduction of thermal information enhances the model’s ability to distinguish objects in complex environments [wang2018rgb, shigematsu2017learning, song2022novel]. Thermal images capture temperature distributions, highlighting object contours, while RGB images deliver abundant color and contextual information that enables a more intuitive understanding of the scene. To leverage inter-modal complementary cues, existing BSOD models can be categorized into early, middle, and late fusion approaches.
@@
Early fusion models typically aggregate bimodal image pairs linearly as input [zhao2020single, 9803225, fu2021siamese]. While this strategy reduces computational load, it fails to prevent the inter-modal noise from interfering with each other. In contrast, late fusion models often employ dual- [pang2020hierarchical, 9454273] or triple-stream branches [ji2021calibrated, cong2022cir] that focus on semantic features or decoder feature fusion. However, these approaches overlook features at early levels that retain important fine-grained spatial information. Therefore, the balanced middle fusion strategy has become the dominant approach in BSOD [song2023potential, zhou2021ccafnet, 9505635, jin2022moadnet, 9926193, 10003255, 10042233, 10127616, wang2024learning]. For instance, Huo et al. [9505635] proposed an ultra- lightweight network with cross-modal filtering and step-wise refinement, while a global illumination module was introduced [9926193] to determine the use of the thermal modality. Zhou et al. [10127616] developed a wavelet-based model that uses knowledge distillation to transfer information from a complex model. Wang et al. [wang2024learning] proposed an adaptive fusion bank, which utilizes basic fusion schemes to address different challenges simultaneously. Despite the significant advances in these models, their performance is constrained by the limited view of convolution operations. Inspired by the attention mechanism employed by the Transformer [vaswani2017attention] which is good at capturing global information, researchers have applied Transformer-based solutions to BSOD [liu2021visual, wu2023transformer, hu2024cross, zhou2023position, 9611276, sun2023catnet, cong2023point, 9869666, 10015667]. Specifically, recent works [hu2024cross, 9611276, sun2023catnet, 9869666] employed Transformer-based backbones to model initial multi-level features. Zhou et al. [zhou2023position] designed an auxiliary task to optimize pixel classification and proposed a Transformer-based decoder. Pang et al. [10015667] proposed view-mixed attention for integrating cross-modal and cross-level features. Cong et al. [cong2023point] introduced position constraints to enhance cross-modal long-range modeling. Although these models effectively address BSOD challenges, they are designed for manually aligned bimodal image pairs and struggle with unaligned instantaneous data. In contrast, our proposed model, AlignSal, is developed for unaligned BSOD data, which can process unaligned RGB and thermal image pairs in real time and exploit the inter-modal complementarity to accurately identify salient objects.
@@
Recently, Tu et al. [tu2022weakly] simulated weakly aligned data by applying affine transformations to aligned bimodal data and used dynamic convolution for weakly aligned BSOD. However, such artificially unaligned data does not fully capture real-world conditions. To address this, Song et al. [10315195] constructed an unaligned BSOD dataset based on AAV views of urban scenes and proposed the MROS model, which performs image alignment followed by fusion. While MROS demonstrates excellent performance, its high computational demands hinder real-time detection capabilities. In order to prioritize both accuracy and real-time performance, we allocate resource-intensive bimodal alignment to the training phase, which shifts part of the computational burden from inference to training. We also design a hierarchical global filtering mechanism with low complexity to handle modal fusion and alignment. Our model achieves competitive performance compared to state-of-the-art models while enabling real-time detection with manageable complexity.
@@ II-B Contrastive Learning (CL)
Contrastive learning (CL) focuses on learning data representations by analyzing the similarities and differences between samples. In CL, the distance between features of similar samples in the embedding space is minimized while that between dissimilar samples is maximized, facilitating the acquisition of meaningful feature representations [hadsell2006dimensionality]. CL has been successfully applied in various fields of computer vision [radford2021learning, wang2021exploring, jiang2024t, sung2024contextrast, wu2023transformer, song2024contrastalign]. For instance, Radford et al. [radford2021learning] proposed a model for aligning visual and textual information through CL, and Jiang et al. [jiang2024t] combined visual and textual prompts within a single model via CL. Song et al. [song2024contrastalign] aligned RGB and radar features using CL to improve the robustness of fusion. Beyond its use for aligning heterogeneous modalities, CL has also been widely studied to enhance inter-pixel discrimination. Wang et al. [wang2021exploring] extended pixel-level CL from a single image to neighboring images. Sung et al. [sung2024contextrast] proposed contextual-based CL to obtain clear feature distinctions. Wu et al. [wu2023transformer] utilized CL to increase the separation between foreground and background pixels in predictions. In this paper, to address the misalignment of RGB and thermal modalities, we leverage CL to align bimodal features at the semantic level and address AAV-based unaligned BSOD.
Figure 2: The overall framework of the proposed AlignSal. The RGB and thermal images are first fed to a dual-stream encoder, which extracts initial bimodal features $\left\{{f_{i}^{r}}\right\}_{i=1}^{4}$ and $\left\{{f_{i}^{t}}\right\}_{i=1}^{4}$. During training, the semantic contrastive alignment loss (SCAL) facilitates the exchange of information between the RGB semantic feature $f_{4}^{r}$ and the thermal semantic feature $f_{4}^{t}$, essentially achieving bimodal alignment. The bimodal features are then aligned and fused at the pixel level by the synchronized alignment fusion (SAF). Finally, a simple decoder integrates the fused features $\left\{{f_{i}^{s}}\right\}_{i=1}^{4}$ from high to low levels, generating the decoder features $\left\{{f_{i}^{d}}\right\}_{i=1}^{4}$ and the final prediction $\mathcal{S}$.
@@ III-A Overview
As shown in Fig. 2, the proposed AlignSal consists of four main components: the dual-stream encoder, the semantic contrastive alignment loss (SCAL), the synchronized alignment fusion (SAF), and the decoder. First, considering that the fast Fourier transform (FFT) offers low complexity while capturing global dependencies, we use the lightweight FFT-based CDFFormer-S18 [tatsunami2024fft] as the encoder. Leveraging the lower inductive bias introduced by FFT [rao2021global, guibas2021adaptive, tatsunami2024fft, 10127616], we employ a Siamese architecture [daudt2018fully, 10127616] for the encoder to extract four levels of RGB features $\left\{{f_{i}^{r}}\right\}_{i=1}^{4}$ and thermal features $\left\{{f_{i}^{t}}\right\}_{i=1}^{4}$, further reducing complexity. During the training phase, the high-level bimodal features $f_{4}^{r}$ and $f_{4}^{t}$, rich in semantic information, are aligned by SCAL to achieve a consistent semantic distribution. Simultaneously, these bimodal features are fed into the corresponding SAF in pairs to achieve pixel-level alignment while capturing inter-modal complementary information. Finally, the outputs from the SAFs $\left\{{f_{i}^{s}}\right\}_{i=1}^{4}$ are passed through a simple decoder to progressively generate up-sampling features $\left\{{f_{i}^{d}}\right\}_{i=1}^{4}$ and the saliency map $\mathcal{S}$.